Solve and check the equation for X
x - x+1 = 8
--- ----- -----
x-2 x x^2-2x
I would first cross out the two x's
Leaving the equation looking like:
0 x+1 = 8
---- - ----- -----
x-2 0 x^2-2x
To go from here I believe I would add x+1= 1x
and subtract x-2 which equals -2x
The equation would look like this:
-2x - 1x = 8
-----
x^2-2x
From here I am lost but I think I became lost a while ago

Answer by checkley71(8403)

X/(X-2)-(X+1)/X=8/(X^2-2X) COMBINE THE FRACTONS ON THE LEFT SIDE OF THE EQUATION & FACTORING THE DENOMINATOR OF THE RIGHT SIDE WE GET
[(X^2)-(X+1)(X-2)]/X(X-2)=8/X(X-2) BECAUSE THE DENOMINATORS ARE EQUAL WE CAN SET THE NUMERATORS EQUAL THUS
X^2-(X^2-X-2)=8
X^2-X^2+X+2=8
X=8-2
X=6 SOLUTION
PROOF
6/(6-2)-(6+1)/6=8/(6^2-2*6)
6/4-7/6=8/(36-12)
6/4-7/6=8/24
6/4-7/6=1/3
(6*6-7*4)/24=1/3
(36-28)/24=1/3
8/24=1/3
1/3=1/3
###############################################
MERRY CHRISTMAS